OpenStudy (anonymous):

solve the system of equations 2x-3y=4 5x+2y=-9

OpenStudy (anonymous):

2x-3y=4 5x+2y=-9 2x=3y+4 x=3/2y +2 5(3/2y+2) +2y=-9 15/2y+10+2Y=-9 9.5y+10=-9 -10 -10 9.5y=-19 /9.5 /9.5 y=-2 2x-3(-2)=4 2x+6=4 -6 -6 2x=-2 x=-1